OpenStudy (anonymous):

Given f(p)=180-0.3p^2, where p is the price in dollars and f(p) is the number of items sold. At what price will maximum revenue be generated?

OpenStudy (anonymous):

revenue= the price* quantity, let's g(p) be a new function, represent revenue \[g(p)=pf(p)=180p-0.3p^3\]

OpenStudy (anonymous):

now just take the derivative of g(p), set it equal to zero, and find the value of p (the price) for maximum revenue as follow:

OpenStudy (anonymous):

\[g'(p)=180-0.9p^2=0 \implies p^2=200 \implies p=\pm10\sqrt2 \] price can never be a negative value, therefore we will take only the positive value, which is \[p=10\sqrt2\] that's the price at which maximum revenue will be generated.

OpenStudy (anonymous):

I hope that makes sense to you

OpenStudy (anonymous):

at p = 0

OpenStudy (anonymous):

^^ p=0 is the price to maximize f(p) clearly, which is the number of sold items. you will never get maximum revenue selling things for free :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

sorry the last step it's \[p=5\sqrt2\] sorry for the typo mistake

OpenStudy (anonymous):

Try www.aceyourcollegeclasses.com

OpenStudy (anonymous):

it is P^2=20 ---> p=5(sqrt2)

OpenStudy (anonymous):

what's wrong with me?? :@ the first answer is right.. lol

OpenStudy (anonymous):

the initial problem is f(p)=180-0.3p^2